OpenStudy (anonymous):

[(30 + 6) − 32] ÷ 9 • 2

OpenStudy (e.cociuba):

[(30 + 6) - 32]/9 * 2 = [(36) - 32]/9 * 2 =4/9 * 2 =?

OpenStudy (e.cociuba):

So when you calculate 4/9 * 2, what do you get?

OpenStudy (anonymous):

is that 4 divided by 9 times 2?

OpenStudy (e.cociuba):

The main thing you want to use is PEMDAS.

OpenStudy (e.cociuba):

Yes it is :)

OpenStudy (anonymous):

[(30 + 6) − 3^2] ÷ 9 • 2 this is actually how the problem should look opps

OpenStudy (anonymous):

Helpp

OpenStudy (e.cociuba):

Ok so your going to solve it exactly like I did on the first one but replace 3^2 as 9 because 3 x 3= 9.

OpenStudy (e.cociuba):

[(30 + 6) − 9] ÷ 9 • 2 [(36) - 9] /9 * 2 = 27/9*2 =?

OpenStudy (e.cociuba):

So, what would you get now?

OpenStudy (anonymous):

6

OpenStudy (e.cociuba):

Correct. Bravo! :)

OpenStudy (e.cociuba):

@Tayloresc do you want me to explain all of this to you? I want to make sure you understand.

OpenStudy (nincompoop):

hmmmmmm are you sure?

OpenStudy (anonymous):

yes

OpenStudy (e.cociuba):

@nincompoop quit confusing him.